OpenStudy (anonymous):

Write an equation for the circle with center (-4, 4) and containing the point (-1, 16). Use the ^ key for the exponents. Write your answer as the example: (x+2)^2+(y-9)^2=100

OpenStudy (anonymous):

(x-4)^2+(y-4)^2=r2 gotta find the radius one sec

OpenStudy (anonymous):

actually it will be \[(x+4)^2+(y-4)^2=153\]

OpenStudy (anonymous):

(x+4)^2 + (y-4)^2 = r^2 radius can be found as folows Sqrt ( 16-4))^2 +( -1+4)^2 = sqrt (144+9) = sqrt 153 r^2 = 153 Answer is (x+4)^2 + (x-4)^2 = 153

OpenStudy (anonymous):

yea thats it, i couldnt remember how to find the radius :oP

OpenStudy (anonymous):

btw you do not need to find the distance, only the square of the distance because that goes on the right hand side. \[(-4+1)^2+(16-4)^2=9+144=153\]

OpenStudy (anonymous):

sorry - the second bracket should be (y-4) not x -4

OpenStudy (anonymous):

maybe you can help me with my question lol

OpenStudy (anonymous):

are u happy with the answers same?

OpenStudy (anonymous):

so its (y-4)? instead of x-4